the marks scored by five students in a test of statistics carrying 100 marks are 50, 60, 50, 60 and 40. A simple random sample of size 4 draws without replacement construct the sampling distribution of sample mean and find the standard error of the sample mean?


1
Expert's answer
2021-07-19T05:47:21-0400

We have population values "50,60,50,60,40," population size "N=5," and sample size "n=4." Thus, the number of possible samples which can be drawn without replacement is



"\dbinom{5}{4}=5"




"mean=\mu=\dfrac{50+60+50+60+40}{5}=52"

"Variance=\sigma^2"

"=\dfrac{1}{5}((50-52)^2+(60-52)^2+(50-52)^2"

"+(60-52)^2+(40-52)^2)=56"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 No & Sample & Mean \\ \hline
 1 & (50, 60, 50, 60) & 55 \\
 \hdashline
 2 & (50, 60, 50, 40) & 50 \\
 \hdashline
3 & (50, 60, 60, 40) & 52.5 \\
 \hdashline
4 & (50, 50, 60, 40) & 50 \\
 \hdashline
5 & (60, 50,60,40) & 52.5 \\
 \hdashline

\end{array}"


The sampling distribution of the sample mean "\bar{x}" and its mean and standard deviation are:


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:}
 \bar{x} & f & f(\bar{x})& \bar{x}f(\bar{x})& \bar{x}^2f(\bar{x}) \\ \hline
 50 & 2 & 2/5 & 20 & 1000 \\
 \hdashline
 52.5 & 2 & 2/5 & 21 & 1102.5 \\
 \hdashline
 55 & 1 & 1/5 & 11 & 605 \\
 \hdashline
 Total & 5 & 1 & 52 & 2707.5 \\
 \hdashline
\end{array}"

"E(\bar{X})=\sum\bar{x}f(\bar{x})=52"

"Var(\bar{X})=\sum\bar{x}^2f(\bar{x})-(\sum\bar{x}f(\bar{x}))^2"

"=2707.5-52^2=3.5"

"E(\bar{X})=52=\mu"

"Var(\bar{X})=3.5=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})=\dfrac{56}{4}(\dfrac{5-4}{5-1})"




"SE=\dfrac{s}{\sqrt{n}}=\dfrac{\sqrt{3.5}}{\sqrt{4}}\approx0.9354"